### v2

# Reinforcement Learning for Control of Evolutionary and Ecological Processes

Reinforcement Learning for Control of Evolutionary and Ecological Processes
Bryce Allen Bagley Mathematical Medicine Group, Stanford University Petritsch Laboratory, Department of Neurosurgery, Stanford Medical School Physician-Scientist Training Program, Stanford Medical School Correspondence: bbagley@stanford.edu Navin Khoshnan Mathematical Medicine Group, Stanford University Petritsch Laboratory, Department of Neurosurgery, Stanford Medical School Claudia K Petritsch Mathematical Medicine Group, Stanford University Petritsch Laboratory, Department of Neurosurgery, Stanford Medical School
SUMMARY
As Evolutionary Dynamics moves from the realm of theory into application, algorithms are needed to move beyond simple models. Yet few such methods exist in the literature. Ecological and physiological factors are known to be central to evolution in realistic contexts, but accounting for them generally renders problems intractable to existing methods. We introduce a formulation of evolutionary games which accounts for ecology and physiology by modeling both as computations and use this to analyze the problem of directed evolution via methods from Reinforcement Learning. This combination enables us to develop first-of-their-kind results on the algorithmic problem of learning to control an evolving population of cells. We prove a complexity bound on eco-evolutionary control in situations with limited prior knowledge of cellular physiology or ecology, give the first results on the most general version of the mathematical problem of directed evolution, and establish a new link between AI and biology.
KEYWORDS
reinforcement learning, evolution, evolutionary dynamics, artificial intelligence, evolutionary game theory, quantitative oncology, mathematical oncology
INTRODUCTION
One of the most productive areas in the study of evolutionary dynamics has been Evolutionary Game Theory (EGT).1, 2, 3, 4, 5, 6, 7 EGT combines elements of the discrete games from economics, the differential games primarily from engineering, and the mathematical and biological literature on evolution and ecology, and uses them to predict how evolutionary processes determine the frequencies of different types in a population. [^1] The term ”types” in EGT is a deliberately general one, as the mathematical tools can be used for the study of genotype frequencies in a population, foraging strategies in macro-organisms, cultural changes, and even collective behavior in response to economic policies.
A key goal of predicting evolutionary dynamics is to be able to intervene and steer evolution towards a desired state. This would represent a quantitative form of directed evolution, an approach of artificial selection already used on bacteria to optimize biomolecular catalysts for industrial manufacturing of pharmacologics, agricultural chemicals, and more.8, 9, 10, 11 This is essentially an approach of applying artificial selection in much the same fashion that natural selection optimizes organisms in nature. While the evolutionary procedures applied in these projects have been relatively simple in comparison with the complexity of evolution in nature, their existing success is indicative of the possibilities which lie in exploiting more of evolutionary – and especially eco-evolutionary – dynamics. Accomplishing this will require mathematical tools for guiding dynamics in a fine-grained manner, and such tools will themselves require the development of new theoretical and computational results. In this paper we expand the theoretical methods available for evolutionary dynamics via integrating contemporary research in control theory and the theory of artificial intelligence.
In medical contexts one finds problems of directed evolution in the choice of therapies for tumors and other pathologies whose ability to evolve drug resistance is a key obstacle in curing many patients. There is been growing interest within clinical medicine regarding the possibility of applying EGT to the development and optimization of cancer therapies, immunology, and infectious disease treatment, but thus far there has been limited ability to integrate the enormous quantities of available biomedical data. Additionally, there are very few studies of treatment courses with meaningful dynamics over time, particularly because of the difficulty of experimenting with combination and time-course therapies. Instead, the current standard of care in cancer is generally to use one or two chemotherapeutics along with surgery and/or radiation. If the patient is not cured, the tumor develops resistance to those drugs, and a new drug or drug combination must be used. However, as discussed in 7, 3, 6, 5 and many others, there is growing evidence that therapies which vary over time in more strategic and complex ways have great potential to improve cancer outcomes. Similarly, bacterial and viral infections develop resistance to antibiotics and antivirals via evolutionary processes very much analogous to those in tumors.
12, 13 A recent resurgence of EGT in the social sciences14, 4 has seen analogous ideas return to substantial use in their intersections with engineering. This is most commonly seen in areas such as transit, energy distribution, computer usage, and other resource problems determined by the behaviors of large numbers of individuals belonging to a smaller number of types which each correspond to some form of pattern in a user’s problem-relevant behavior.15, 16, 17 Our work is distinct from this approach in that we are not using heuristic genetic algorithms for control, but rather are using rigorous methods for the control of evolution.
In economics, engineering, and biomedical research alike, here has been relatively little work applying contemporary advancements in control theory and related disciplines to questions of evolutionary systems. There is a recent discussion of the use of dynamic programming for the optimization of cancer therapy,18 however, we have been unable to find any such work beyond that paper. The massive advancements in the mathematical study of control systems and artificial intelligence from recent decades have not yet crossed over into the study of evolutionary dynamics. This represents a fertile ground for the use of new tools and the answering of previously unstudied – and in some cases perhaps even previously unposable – questions.
Here we explore two key ideas. First, questions of how to prove a rigorous algorithmic bound on the computational problem of optimal control of evolution. Second, the duality in the nature of the underlying process as both a competition and a computation. The former has clear connections with what we discuss above about the history and present state of the cross-talk between game theory and evolutionary biology. [^2] This conceptual view is ubiquitous in theoretical biophysics, systems biology, and adjacent fields, and a recent review article offers a helpful explanation of them in terms of potential therapeutic applications.19 The latter, however, is perhaps a subtler one, and is related to a radical yet compelling proposal by Krakauer, Bertschinger, Olbrich, Flack, and Ay – theoretical biophysicists and theoretical biologists working in the space of complex systems science.20 They propose that individuality exists not as a binary property, but a spectral one with its origins in information theory. Instead of arbitrary human-imposed boundaries, one instead looks at how the mutual information of different partitionings varies and thus defines degrees of individuality.
We do not make use of the explicit mathematical description they offer, but take inspiration from its conceptual implications. Instead, we note that from a pure information perspective a change in a molecule’s structure is fundamentally no more or less an information transform than a molecule being moved from one partition to another. We can imagine attaching to any molecule a hypothetical string of bits to serve as an address of sorts. One value for the bit string could say that it is within a biological cell, another that it is outside the cell. This equivalence will be discussed further in Computational Duality and Multicules of the .
Further, we make use of Osband and Van Roy’s 21, 22 work on model-based reinforcement learning (RL) of Markov Decision Processes, in which they proved the first unified regret bounds for model-based RL of quadratic systems – like EGT – which to our knowledge remain state of the art to this day. Linking these results from the theory of artificial intelligence with the specific function classes [^3] A term used in machine learning and artificial intelligence to denote a collection of functions with one or more relevant properties in common. of interest in EGT, we introduce and develop a mathematical formalism for an evolutionary control theory via a combination of EGT, reinforcement learning, and control systems. In addition to defining and exploring this formalism, we derive bounds on learning to control directed evolution which match the state-of-the-art bounds in the theory of artificial intelligence more broadly. However, much of the value of theory comes from eventual application to practice, and as discussed by Traulsen and Glynatsi there is a growing divide between theory and practice in EGT.23 Our goal in this paper is to develop a novel approach which introduces new techniques into the theoretical literature of EGT while focusing on its ultimate use in experiment and application.
In the standard formulation of EGT,1, 2, 4 one studies a vector $c$ of $q$ types – $c_{1},c_{2},...,c_{q}$ – which interact within some environment with the goal of predicting what fraction of the total population each type will comprise over time. The interactions between types are modeled as an interaction matrix $\mathbf{F}$ such that $f_{ij}$ predicts how the $j$-th type will impact the $i$-th type. The unadjusted fitness of a type $c_{i}$ – that is, the rate at which its population will grow – is thus defined by $\mathbf{F}_{i}c$, where $\mathbf{F}_{i}$ is the $i$-th row of $\mathbf{F}$. Because we are concerned with what fractions of the population each type comprises, these fitnesses must be normalized. Thus, the average fitness of all types is computed via weighting by population fraction, and this is subtracted from the fitnesses of all types to normalize population growth.
While in reality populations essentially never remain static in total size, this fractional representation is far more useful for mathematical analysis and in general does not diminish the value of the analysis. The growth of a type’s sub-population is proportional to its current size, so the total formulation of a basic evolutionary game theory problem is as follows.
For a specific type,
[…]
Extending to all types present, we have
[…]
Here $\mathbf{C}_{a}$ is a diagonal matrix whose entries are the elements of $c$, defined explicitly as
$\mathbf{C}_{a}=\sum_{i=1}^{q}\mathbf{E}_{i}c^{T}\vec{e}_{i}$ (2)
Where here $\vec{e}_{i}$ is the vector with all zero entries save for a $1$ at entry $i$, and $\mathbf{E}_{i}$ is the $q\times q$ matrix with the $q$-th main diagonal entry as 1 and all other entries as 0. We have subscripted this matrix $\mathbf{C}_{a}$ for convenience and cleaner notation in certain more complex calculations to come.
Additionally, $c^{\hskip 2.84526ptT}\hskip 2.84526pt\mathbf{F}c$ is the scalar representing the population-fraction weighted sum of raw fitnesses of all types. This can be more understood as first computing the fitness of each type and then taking the weighted average for the population by multiplying each individual fitness by its population fraction and summing.
This general form of nonlinear dynamical system can thus be used to predict how the composition of the overall population will change over time. Extensions often account for mutations, which in the continuous case essentially transfer infinitesimal fractions of the population from one type to another, but the general principle holds.1, 2, 4, 24
If mutations are to be incorporated, this would be done via the addition of a linear mapping $\mathbf{M}$ which describes the average rate at which each respective type mutates into any others. This mapping would be a symmetric matrix with diagonal entries $1-\delta$, with the off-diagonal entries representing these mutations being some distribution which sums to $\delta$.
In the late 20th century (and especially the 21st thus far) quantitative biology largely split into theoretical biology and bioinformatics. Compared with the deeply interlinked nature of theory and data analysis in the first half of the 20th century, comparatively little interaction occurred between them in the latter half. Despite being theoretical, EGT has still provided useful insights even in abstract yet practical contexts. It offers explanations of why intuitively strange situations in ecology can remain stable, and in recent years has seen early application to analyzing tumor progression and microbial evolution of antibiotic resistance. However, these efforts have in some ways been constrained by the methods used. In most applications the number of types $q$ might be three or four, and the entries of $\mathbf{F}$ will be bird’s-eye-view estimates from large-scale observations of bacterial and cancer cell populations.
Bioinformatics tools which have emerged in recent decades may allow quantitative researchers to generate higher-fidelity approximations of $\mathbf{F}$ for $q$ large enough to make clinically relevant predictions beyond the simple test cases being assessed at present. Such work is critical, and a goal of this study is to begin the creation of a new class of tools for such researchers to use. In particular, work in the area of network inference algorithms has made it increasingly possible to estimate interaction effects within cells and between them. This represents an opportunity for a new approach to eco-evolutionary dynamics which can make use both of contemporary methods in biomedical data sciences and of rich theory.
RESULTS
As discussed in the Methods subsection Computational Duality and Multicules, one of the new methods we introduce is to model multiple types of biophysical processes as all representing a single unifying sort of computation. For example, the conversion of one molecule into another by an enzymatic process is a computation as it changes the set of interactions the matter from the original molecule can engage in. Likewise, the movement of a molecule from a cell into the intercellular space – or vice versa – also changes the set of interactions that molecule can engage in. We introduce the term multicules to denote the units of this unifying computational representation.
There are a few options in terms of how to model a system where data is available in the form of values for $c$ and limited data on $x$, while the dynamics themselves are in truth defined in terms of $x$. One option is to assume a given estimate of $c$ maps onto an estimated $x$ (and vice versa), and have the model be a hybrid of dynamics over $c$ and $x$. For the purposes of this study, the approach will be to assume a pseudo-invertable [^4] See Section 1 for a discussion of the validity of the assumption of pseudo-invertability. linear map $\mathbf{B}x=c$ which is time-invariant. Incorporating noise will be relevant later, so we will use the notation $\mathbf{\tilde{B}}=\mathbf{B}x+\xi$ with $\xi$ being a vector of zero-mean noise. For the time being, however, we will assume perfect ”estimates” and ignore the $\xi$ component of $\mathbf{\tilde{B}}$. Further, we define a vector of inputs to the ecology $u$, and a matrix $\mathbf{Z}$ mapping inputs to multicules.
However, it is not in fact required that we have any knowledge of $\mathbf{B}$, $\mathbf{G}$, or any other parameter matrices a priori. This is where the application of reinforcement learning will become central to later sections of the study. For now, we must first define a model of the eco-evolutionary process. The rest of this section will be a discussion of the model and how we can convert it into a Partially Observable Markov Decision Process (POMDP) so as to enable the use of the extensive research in control systems engineering and the theory of artificial intelligence.
0.1 Dynamical System Model for Multicule Eco-Evolutionary Dynamics
The underlying mathematical model is described entirely with respect to multicule ”computations” on $x$ as
$d_{t}c=\mathbf{M}\mathbf{C_{B}}^{T}\left(\mathbf{G}x-\vec{\hskip 1.42262pt1}(\mathbf{B}x)^{T}(\mathbf{G}x)\right)+\tilde{\mathbf{Z}}u$ (3)
With $\mathbf{C}_{B}^{T}$ being a matrix whose diagonals are the entries of $\mathbf{B}x$. In other words, $\mathbf{C}_{B}$ is defined as
$\mathbf{C}_{B}^{T}=\sum_{i=1}^{q}\mathbf{E}_{i}(\mathbf{B}x)^{T}\vec{e}_{i}$ (4)
Our definition of $\mathbf{B}$ is such that this would be equivalent to the earlier $\mathbf{C}_{a}$ if our inferences were perfect, but this would require a perfect linear model of all molecular interactions involved – a fundamental impossibility given the nonlinearity of biological processes. Additionally, values of $c$ – and thus $\mathbf{C}_{a}$ – would be noisy estimates from data.
We note a potentially non-obvious feature of Equation 3. We refer to this model as being a quadratic dynamical system, yet it is clear that the term $\vec{1}(\mathbf{B}x)^{T}(\mathbf{G}x)$ is quadratic but will then be multiplied by $\mathbf{C}_{B}$ – which is a function of $x$. However, because of the structure of the problem – as seen in the fact that we have a scalar value multiplied by $\vec{1}$ so that it will be applied to all vector elements of $d_{t}c$ – this is equivalent to defining an additional variable. This would be a book-keeping variable, in some sense, such that we have $\vec{1}(\mathbf{B}x)^{T}(\mathbf{G}x)=\vec{1}x_{\text{book}}$ separate from the actual $x$ entries. Its dynamics would then be $d_{t}x_{\text{book}}=(\mathbf{B}x)^{T}(\mathbf{G}x)$ to preserve system structure. As a result, every component of the dynamical system is either of either linear or quadratic order. The same approach can easily be extended to the full model of Equation 8 and its equivalent 2nd-order equation.
We assume dynamics to alterations in the ecology can be approximated as some multidimensional impulse at time $t_{0}$ which decays back to a baseline. $\mathbf{U}$ is a diagonal matrix analogous to $\mathbf{C}_{B}$, but for the vector of inputs $u$ as
$\mathbf{U}=\sum_{i=1}^{m}\mathbf{E}_{i}u^{T}\vec{e}_{i}$ (5)
Other functional forms are of course possible, and in a wide variety of cases we anticipate that our proofs will apply equally to other models. In this study we will only analyze this particular form of ecological model because the ability to construct high-dimensional sigmoids from it allows the model to describe a broad range of ecological processes. This sigmoid model is introduced below.
0.2 Non-static Ecology
Additionally, we would like inputs to be able to modify the ecology of the system and physiology of the organisms within it in ways not directly captured by a mapping from $u$ to $x$. To do this, we define
$\mathbf{G}(t)=\mathbf{G}_{0}+\mathbf{\Delta}\mathbf{U}\mathbf{V}$ (6)
It is important to note that the form of $\mathbf{\Delta}\mathbf{U}\mathbf{V}$ captures modifications to the ecology of an eco-evolutionary system which are separate from the direct multicules introduced via an external input into the system. The matrices $\mathbf{\Delta}$ and $\mathbf{V}$ are needed to reshape the dynamics of $\mathbf{U}$, and the three matrices in concert describe the effects of inputs on the ecological state of the model just as $\mathbf{Z}$ describes the inputs’ direct mapping to $x$ where relevant.
To capture the decay process post-impulse, we use a diagonal matrix of decay rates $\Xi$ modifying $\mathbf{\Delta}\mathbf{U}\mathbf{V}$ into $\mathbf{\Delta}\tilde{\mathbf{U}}e^{-\Xi t}\mathbf{V}$, and have
$d_{t}\mathbf{G}=-\mathbf{\Delta}\mathbf{\tilde{U}}\mathbf{\Xi}e^{\mathbf{-\mathbf{\Xi}}t}\mathbf{V}$ (7)
One can model a shift from one baseline ecology to another simply by having the midpoint between the two baselines be analogous to one impulse decaying backwards in time towards an original $\mathbf{G}_{0}$, and a second impulse as meeting in the middle with respect to some new baseline $\mathbf{G}_{0}^{\prime}$.
If we proceed through the matrix multiplications as discussed in Section 1, we derive the completed model of the evolutionary dynamics with the estimation procedures taken into account.
1 d_t x &= B^-1C_B(MGx - →1(x^TB^TGx)) + Zu
d_t G = -Δ~UΞe^tV
Note that $\mathbf{Z}$ is equivalent to $\mathbf{B}^{-1}\tilde{\mathbf{Z}}$, to streamline notation. Entries of $\mathbf{Z}$ serve to map external inputs into the multicules they become within the system, in this study according to the functional form of $u$ as decays from a given input point. In terms of pharmaceuticals, the term including $\mathbf{U}$ can be thought of as the effects of active drug molecules, and $\mathbf{Z}u$ as combining the rates of excretion and the rates of metabolism into pharmacologically inactive compounds.
0.3 POMDP for Multicule Eco-evolutionary Models
This is a nonlinear dynamical system just as with Equation 1, and in the next section we proceed to a discussion of how model predictive control can be used to influence it. Finally, for the sake of our later analysis, it can be useful to convert the system in Equation 8 into a single, equivalent higher-order equation. The derivation is somewhat messy, so they can be found in Section 1 as well. Recalling our earlier mention of a ”bookkeeping” variable, we note that it is not illustrative in the context of the model’s dynamics, but does verify the model’s property of being a quadratic dynamical system. That property will be important later, as regret bounds for model-predictive reinforcement learning are not known for all but a small number of very specific equations and a much smaller number of model classes – linear, quadratic, and generalized linear models being the only such general classes for which there are to our knowledge adequate bounds.
To condense our notation for the proceeding sections, we will discretize the 2nd-order dynamical model into a Markov Decision Process (MDP) transfer function.
$\Psi(x^{\prime}|x,u)=\psi(x,u)+\vec{\xi}(x_{k+1}-x_{k})\cdot(\delta t)=\frac{d^{2}x}{dt^{2}}(\delta t)^{2}+\vec{\xi}(x_{k+1}-x_{k})\cdot(\delta t)$ (9)
with $\vec{\xi}$ a zero-mean sub-Gaussian noise vector to account for model error (and absorbing the earlier $\xi$ term), multiplied by $(\delta t)$ so that it is proportional to the time-step length. There is thus a parameter $\sigma_{\psi}$ describing the maximum variance across entries of this sub-Gaussian noise vector. The noise is sub-Gaussian rather than Gaussian to prevent the mathematical difficulties connected with the MDP RL analysis on an unbounded range of outcomes, and this is also physically correct given conservation of mass. Because we can only observe a subset of the system’s components at any given time, this MDP is termed ”Partially Observable” and is thus a POMDP.
As we will soon show, the use of reinforcement learning means that there are minimal requirements of a-priori knowledge of a system. Beyond specifying definitions of multicules and input compounds, we can leave all parameter matrices unknown at the start of the problem. this will of course increase the number of trials needed to learn a model when compared with learning given partial prior knowledge of parameters, but the worst-case bounds described in LABEL:accuracy_section are valid for the cases of any amount of prior knowledge, including the case of no such knowledge. Armed with this single-equation POMDP, we proceed to develop the optimal control results.
0.4 Parameter Learning
Summarizing the evolutionary game using the time-evolution operator $\Psi(x(k+1)|x(k),u(k))$ from Equation 27, describing how the system changes based on its present state and inputs at some time $k$, we can construct a Bellman function for it. The full discussion can be found in Section 2 of the Supplemental Information, but to summarize we have the following:
$V(x(k),u(k))=L(x(\tau),u(\tau))+\gamma\int_{X}dx\Psi(x(k+1)|x(k),u(k))V(x(k+1),u(k+1)$ (10)
Here $V(\cdot,\cdot)$ is the value at a given time based on the state and input, and $L(\cdot,\cdot)$ is the reward associated with the final state which would result at time $\tau$ following a given input sequence $\{u(k),u(k+1),...,u(\tau)\}$ given the current state $x(k)$.
The remaining key component of our derivations is the learning of parameter matrices themselves. Beginning with some set of matrices inferred from experiments, the expectation values and uncertainty in those inferences define respective prior distributions over the parameter matrices. Here we are agnostic as to the method used to generate these prior distributions.
One will inevitably need to select a subset of variables to create a Reduced-Order Model (ROM) for the POMDP $\Psi$, given the enormous dimensionality of the ideal approximation: a Full-Order Model incorporating all multicules involved in the system. There will be a lower bound on the error resulting from the difference between the ROM and the FOM, expressed as the residual, which is the difference between the state resulting from a ROM and that which would be seen for the same variable subset in an FOM. We denote the magnitude of this error as $\varepsilon_{O}$. A number of authors have described error bounds for various types of dynamical systems, and this continues to be an area of active research.25, 26, 27, 28, 29, 30 To our knowledge the best bound on this error for nonlinear dynamical systems is given by Chellappa et al in 31. Their method is rather technical, and translating their error bounds to the context of our problem in detail is beyond the scope of this paper. We note the existence of such methods both to acknowledge an interesting and highly relevant body of literature, and to indicate the direction one could pursue in proving detailed bounds on $\varepsilon_{O}$. In addition, there is some fundamental error limit $\varepsilon_{\text{min}}$ below which a quadratic model for a given biological system could not be further optimized, regardless of whether or not it were reduced-order. Thus the problem we analyze here is more precisely one of reinforcement learning to minimize any error beyond the minimum set by those two limits, and optimal treatment selection given such an optimized model.
The error bounds for reinforcement learning of POMDPs proven across 32, 22, 21 are both state of the art [^5] Per our literature review and personal communication with B. Van Roy and M. Kochenderfer. and highly general compared with most results in the field, which makes them very suitable for this problem. A lengthier discussion of these bounds in our case can be found in Supplemental Information Section 4, and relate specifically to the algorithm Posterior Sampling Reinforcement Learning (PSRL).22 In summary, we use their results to derive regret for using the PSRL algorithm to both learn the parameter matrices and select inputs to optimize trajectories through evolutionary space.
Because $\varepsilon_{O}$ is separate from the $\xi$ term which we would like to minimize – the goal of using RL to improve our model – there will be a baseline accumulation of error in addition to $\varepsilon_{k}$, the error from the estimation of the integral portion of the Bellman operator. The derivation of the function describing $\varepsilon_{k}$ is discussed in Section 3 of the Supplemental Information. They will be bounded by the sum of their absolute values, so we know that the error from our approximations (as opposed to the error from sub-optimal parameters, which PSRL is used to minimize) will be bounded as $\varepsilon\leq|\varepsilon_{k}|+|\varepsilon_{O}|+|\varepsilon_{\text{min}}|$, as this would be the case if there were only constructive interference among the error values.
We use $\{\epsilon\}$ to denote the triplet of upper bounds on the errors. First $\varepsilon$, the upper bound on unavoidable error. Second and third are $\epsilon_{R}$ and $\epsilon_{\psi}$, the errors between the inferred and actual functions for each of reward and transition process. $\epsilon_{R}$ corresponds to learning the $\vec{\omega}$ for the reward function if it is not fully specified in advance, and likewise $\epsilon_{\psi}$ for learning the model $\Psi$.
$\mathbb{E}[\tilde{\mathbf{reg}}](T,\Psi,R,\{\epsilon\})=\tilde{\mathcal{O}}\Bigg{(}$ $\sqrt{T}\Bigg{[}\sigma_{R}(m+n)\sqrt{\log\left[\frac{\|\vec{\omega}\|_{2}\|\{x,u\}\|_{\max}}{\epsilon_{R}}\right]}$ (11)
$\hskip 5.69054pt+\mathbb{E}[\mathcal{K}]\sigma_{\psi}n\sqrt{\log\left[\frac{n\|\mathbf{A}\|_{1}\|\mathbf{A}\|_{\infty}\|x\|_{\max}}{\epsilon_{\psi}}\right]}\Bigg{]}\Bigg{)}$
$\mathbf{A}$ is a combined model parameter matrix being learned, [^6] See Supplemental Information Section 1 for a discussion of how the linear quadratic model of $\Psi$ can be converted into a single quadratic function. and $\tilde{\mathbf{reg}}$ is the regret over all learning episodes. Because there are multiple ways of formulating the problem without altering these bounds, as discussed in Supplemental Information Section 1, we do not specify $\mathbf{A}$ in terms of $\mathbf{\Delta}$, $\mathbf{V}$, $\mathbf{\Xi}$, $\mathbf{B}$, $\mathbf{G}$. In each case, it will still hold that $\mathbf{A}\in\mathbb{R}^{\mathcal{O}(n)\times\mathcal{O}(n)}$, so it does not impact our analysis at this level of granularity. Here $\|\mathbf{A}\|_{1}$ and $\|\mathbf{A}\|_{\infty}$ are respectively the maximum column-wise and row-wise sums in the matrix, $\|\mathbf{A}\|_{2}\leq\sqrt{\|\mathbf{A}\|_{1}\|\mathbf{A}\|_{\infty}}$ is the induced $\mathcal{L}_{2}$ norm of the matrix, equivalently its largest singular value, and $n$ is once again the dimension of the multicule vector.
One may notice that despite our discussion of $\varepsilon$ and its components $\varepsilon_{0}$, $\varepsilon_{\text{min}}$, and $\varepsilon_{k}$, the variable $\varepsilon$ does not appear in the final bound. This is because the expectation of its contribution to the regret is in fact 00, with small variance on the scale of the state vector lengths we anticipate this algorithm being applied to. A discussion of this and proof of the distributional properties is included in Supplemental Information Section 5.
The sparseness of $\mathbf{G}$, $\mathbf{U}$, and others due to the rather tight bounds which biology itself imposes upon the matrices combine to further improve the bound beyond what we show in an asymptotic bound. That is, multicules are in general only directly convertible into a small number of other multicules, and the impacts of other processes not directly modeled can reasonably be assumed to be distributed in a Gaussian fashion by application of the Central Limit Theorem. Additionally, it is reasonable to assume that most aspects of physiology will not change drastically over time on average. As a consequence of these factors, for most $\mathbf{A}$ of interest in this problem, $\|\mathbf{A}\|_{1}$ and $\|\mathbf{A}\|_{\infty}$ will likely be $<<n\delta t$.
The portion of total regret accessible to learning grows as the $\sqrt{T}$, but this sublinear increase occurs across linear time and so the regret per unit time decays as $\frac{1}{\sqrt{T}}$, meaning that our learning converges. While the unlearnable component of regret grows linearly, it is zero mean and has a variance which converges to zero for large $n$, yielding an effective overall bound.
0.5 Computational Complexity
We show in Section 1 of the Supplemental Information that the time complexity for model evaluation is $\mathcal{O}(qn^{2})$, and the number of evaluations per timestep for a given degree of error in terms of integral quadrature sample count is calculated in Section 3 of the SI to be $\varepsilon_{k}(s)=\mathcal{O}\left(n^{2}\exp\left(\frac{-s^{2}}{n^{2}}\right)\right)$.
Supposing we want a final error bound of some specific value $\epsilon_{f}=\inf\{\epsilon_{R},\epsilon_{\psi}\}$, each round defining a new $k_{0}$ and requiring assessment across the $\tau-k_{0}$ planning range, and require a quadrature sample count for each time-step sufficient to reach $\epsilon_{f}$. We compute the sample number requirement per time step by inverting the asymptotic bound
$s=\mathcal{O}\left(n\sqrt{\ln\left(\frac{n}{\epsilon_{f}}\right)}\right)$
This gives samples per timestep analyzed, and given the $\mathcal{O}(s^{3})$ complexity of FWLSBQ the total complexity for error $\epsilon_{f}$ is
$\mathcal{O}=\left(n^{6}q^{3}\cdot\hskip 2.84526pt^{3}\sqrt{\ln\left(\frac{n}{\epsilon_{f}}\right)}\right)$
This bound at first appears concerning given the $n^{6}$ term, but because of the practical limitations of experiments we expect $n$ to be on the order of hundreds at most. Additionally, the exponential convergence of FWLSBQ, and the physical limits on a useful $\epsilon_{f}$ imposed by the inevitable level of noise involved in any biological system further diminish the likely computational demands in practice. We know that for each measurement of the system (occurring at intervals of t) we evaluate over the planning range of $\tau$ future timesteps. So the asymptotic complexity is then the product of the total operational time $T$, the planning range length $\tau$, and the quadrature complexity per timestep in planning. Thus the method as a whole has time complexity
[…]
Given the speed limitations of biological processes, we anticipate the timescales available for computational analysis mean that with appropriate refinement for implementation the bounds should make this approach feasible for near-term application.
DISCUSSION
Though a discussion of the practical learning procedure itself is beyond the scope of this study given it will likely be a contextual decision based on the desired application of directed evolution (clinical, industrial, scientific, etc.), our theoretical analysis stands on its own as a first-of-its kind approach for addressing one of the central questions of applied evolutionary theory. The interoperability of Computational duality makes it particularly suitable for the monitoring and course-correction of evolutionary processes, given the multicules corresponding to an intercellular compartment provide a ready readout for the state of the system provided there are adequately accurate inference methods for translating back and forth with experimental data. We expect such inference methods to be very feasible.
We see a number of promising directions for future work, including 1) extending the results to nonlinear transforms $c=\mathbf{B}(x)$ rather than the linear one we assume, 2) analyzing the potential impacts of the difference between the timescale of the respective dynamics of $x$ and $u$, 3) tightening the bounds on quadrature estimate by accounting for sparcity in parameter matrices, and of course 5) applying our approach or variants of it to experimental problems.
In summary, this study represents three key advancements. First, a novel mathematical formulation of Computational duality for biophysical dynamics which seamlessly links ecology and evolution into a single unified process – just as is of course the reality in nature. Second, Computational duality’s universality and ability to more richly interface with the mechanistic properties of biosystems has the potential to strengthen the link between theory and experiment in further highly novel ways. Third, by establishing for the first time the utility of tools from modern artificial intelligence theory cross-applied to the mathematical study of evolution and the biophysical questions therein, we find that in addition to the established relationship of neuroscience and AI this offers a new explicit link between biophysical processes and that of artificial intelligence. All three represent interesting directions for future work of a highly interdisciplinary nature.
Limitations of the study
Because we sought to provide a model that could interact directly with experiments, certain decisions were made which increase the conceptual and computational complexity of the problem. For example, the mathematical model itself is notably more complex than those typically used in evolutionary dynamics. However, each of the modeling decisions was carefully made so that as much prior experimental data as possible could be incorporated into the model before any training might begin. Our goal was to strike a balance between the capacity for integration with physiological experiments – where simpler models have often struggled – and the need to retain a model tractable for algorithmic analysis so that rigorous statements could be made about an interpretable model. Proving bounds for PSRL is straightforward for the coarse-grained models typical in evolutionary dynamics, so in a sense that result is incorporated here as a limit case of our work.
Because the values of $\varepsilon_{O}$ and $\varepsilon_{\text{min}}$ are difficult to estimate in the absence of experimental data, this represents a key aspect which we do not presently know of any means of analyzing mathematically a priori. Methods such as those described in the above discussion of Reduced-Order Models may offer useful tools, but will require substantial calculations in order to prove bounds on $\varepsilon_{O}$. By contrast, $\varepsilon_{\text{min}}$ will almost certainly need to be determined experimentally for each evolving system which one might wish to study. Because of this, it is entirely possible and perhaps inevitable that our approach and derivatives of it will be useful for certain evolving systems, but not prove effective for others.
Ultimately, this paper seeks to be a first step towards incorporating contemporary reinforcement learning with evolutionary dynamics. Significant additional empirical and theoretical work remains, but this provides a foundation for such investigation.
METHODS
Problem Statement
Evolutionary game theoretic models are necessarily nonlinear dynamical systems, and in practice at present it is generally only possible to take estimates of their states occasionally due to the need for genetic sequencing or molecular testing of some sort. Thus, any effort to direct evolution will require the application of digital control.33, 34 In reality, $\mathbf{F}$ is a phenomenological representation of some underlying nonlinear system describing the dynamics of biochemistry and signalling processes within and between biological cells and organisms. Recent improvements in biological data-gathering technology 35, 36 enable the inference of a linear approximation $\mathbf{G}x$ of an underlying nonlinear system for an $n_{0}$-dimensional vector $x$ of molecular levels within cells. It should be noted that $\mathbf{G}$ will be a sparse matrix, as any given molecule will typically only interact meaningfully with a small number of other molecules. Exceptions absolutely exist, such as the central importance of glucose [^7] A sugar which serves as the body’s main source of energy system-wide. and ATP [^8] A high-energy molecule synthesized from a variety of energy sources, which is used to fuel a wide array of cellular processes., but by and large $\mathbf{G}$ will be quite sparse. [^9] That is, most of the elements of the matrix will be zero.
The complexity of biological systems presents substantial challenges, but this complexity also produces enormous opportunities if methods can be developed for deliberately and precisely controlling them.37, 38, 36 Given there are modeling principles which it is believed should transfer over to complex, data-driven, real-world cases, Model Predictive Control appears to be the most viable option for solving this complex MIMO [^10] Multiple Input, Multiple Output control problem.39, 40
Assuming such a $\mathbf{G}$ as described above, the question will be as follows. Given the ability to modify some comparatively small subset of individual entries of $\mathbf{G}$ in the form $\mathbf{G}_{ij}\rightarrow\mathbf{\hat{G}}_{ij}$ such that $\mathbf{G}_{ij}-\mathbf{\hat{G}}_{ij}$ is sufficiently small for a given application, [^11] We do not precisely define sufficiently small precisely because we anticipate it could vary significantly from application to application. the goal is to develop algorithms for predicting and controlling the evolutionary system. In particular, to make predictions of changes produced by targeted selections of the indices for nonzero $\mathbf{G}_{ij}-\mathbf{\hat{G}}_{ij}$.
This yields the equation
$d_{t}c=\mathbf{M}\mathbf{C}_{a}\left(\mathbf{G}x-\vec{1}c^{\hskip 2.84526ptT}(\mathbf{G}x)\right)$ (13)
where we have in essence simply replaced $\mathbf{F}c$ with $\mathbf{G}x$ as per the discussion above, and $\mathbf{M}$ is a mapping describing the average rate of mutations converting one cell type into another.
In the context of reinforcement learning, policies are sets of control signals which are input into a system. Given an specific objective and system there will be one or more optimal policies and many sub-optimal ones generally across a spectrum of quality. This objective is termed the ”reward” or ”value” function, and is defined for a specific problem. Finally, ”regret” is used as a metric of the efficacy of some algorithm or approach. Put simply, regret – typically considered as an expectation value – is the difference between the ”reward” obtained via a policy and the optimal reward that could have been obtained via a perfect policy.
For the problem of directed evolution, the value function will in general be a function of the various cell types being considered, with more desirable types having increasingly positive (or less negative) contributions to the reward. The vector of these cell types’ respective frequencies at time $k$ is represented by $x(k)$. However, experimental supplies are not free, and in a medical context treatments carry with them side effects so an additional term will incorporate their costs. Inputs are represented as $u(k)$.
For mathematical tractability, we use a linear reward function which at time $k$ for state $x$ and input $u$ is
$R(x(k),u(k))=\vec{\omega}^{T}x(k)-\vec{\mu}^{T}u(k)$
Specific $\vec{\omega}$ and $\vec{\mu}$ are determined by the goals in some problem of interest. In this study these goals are set by the humans aiming to control some eco-evolutionary process, but extensions to this model could define it based on an ecology itself to study natural evolution. [^12] For instance, one could set $u(k)$ to zero, use a different ecology function than we do here, and have $R(x(k))$ be simply maximization over replication rates. In other words, survival of the fittest.
There are a few options in terms of how to model a system where data is available in the form of values for $c$, while the dynamics themselves are in truth defined in terms of $x$. One option is to assume a given estimate of $c$ maps onto an estimated $x$ (and vice versa), and have the model be a hybrid of dynamics over $c$ and $x$. For the purposes of this study, the approach will be to assume a pseudo-invertable [^13] See Section 1 for a discussion of the validity of the assumption of pseudo-invertability. linear map $\mathbf{B}x=c$ which is time-invariant. Incorporating noise will be relevant later, so we will use the notation $\mathbf{\tilde{B}}=\mathbf{B}x+\xi$ with $\xi$ being a vector of zero-mean noise. For the time being, however, we will assume perfect ”estimates” and ignore the $\xi$ component of $\mathbf{\tilde{B}}$. Further, we define a vector of inputs to the ecology $u$, and a matrix $\mathbf{Z}$ mapping inputs to multicules.
However, it is not in fact required that we have any knowledge of $\mathbf{B}$, $\mathbf{G}$, or any other parameter matrices a priori. This is where the application of reinforcement learning will become central to later sections of the study. For now, we must first define a model of the eco-evolutionary process. The rest of this section will be a discussion of the model and how we can convert it into a Partially Observable Markov Decision Process (POMDP) so as to enable the use of the extensive research in control systems engineering and the theory of artificial intelligence.
Computational Duality and Multicules
As mentioned in the introduction, the complexity of eco-evolutionary processes can be enormous. Tracking the positions of even a negligible fraction of molecules involved in such processes is not only infeasible but almost certainly unproductive. However, there is value in being able to capture key features of spatial dynamics as well as more direct biochemical and evolutionary ones. One simplified model which offers a good balance between the two is to analyze the spatial component of a system as a set of discrete partitions.
We can imagine attaching to any molecule a hypothetical string of bits to serve as an address of sorts. One value for the bit string could say that it is within a biological cell, another that it is outside the cell. If more fidelity is desired, one could denote intranuclear, [^14] That is, within the nucleus of a cell. cytosolic, [^15] Within the region between a cell’s nucleus and its outer boundary. or extrecellular. Alternatively (or going further), one could create separate label string values for any given partition of interest, such as one value for the shared space between a group of cells and then additionally one value each for every type of cell as described in LABEL:section_background_defining_egt.
In this framework there will in general be no fundamental difference between a molecule moving between partitions vs. being chemically transformed as it undergoes the same motion. That is to say, a process which moves a molecule $\text{mol}_{1}$ from partition $\text{Par}_{1}$ to partition $\text{Par}_{2}$ is mathematically analogous to another process which took $\text{mol}_{1}$ and converted it into $\text{mol}_{2}$ as it is moved from $\text{Par}_{1}$ to $\text{Par}_{2}$. These two processes are dual to each other, in the sense in which physicists use the term duality. They are equivalent descriptions of the same underlying process. The fact that this is not the physical reality of the situation should in general not hamper its efficacy as a modeling tool, and from an epistemological perspective that is likely all any duality from physics could claim to be as well.
Using $\text{mol}_{i}$ to denote molecules and $\text{Par}_{j}$ to denote partitions, [^16] With $i$ and $j$ arbitrary indices and not related to any other use of the same index variables elsewhere in the paper, the equivalence of $\{\text{mol}_{1},\text{Par}_{1}\}\rightarrow\{\text{mol}_{1},\text{Par}_{2}\}\equiv\{\text{mol}_{1},\text{Par}_{1}\}\rightarrow\{\text{mol}_{2},\text{Par}_{2}\}$ requires introducing language for the relationship between $\{\text{mol}_{1},\text{Par}_{2}\}$ and $\{\text{mol}_{2},\text{Par}_{2}\}$. We might term these partition-distinguished molecules particules, but will instead use the term multicules on account of the extensions beyond partitions as additional properties – and to avoid confusion with the particles of physics. We term the duality itself Computational duality.
Computational duality allows us to capture richer spatial aspects of eco-evolutionary dynamics in the molecular reformulation of EGT discussed above, as we can account for molecule placement simply by creating multiple entries in $x$ which correspond to the same physical molecular structure across as many partitions as relevant for a given problem. For $q$ cell types of interest, one extracellular compartment, and $n_{0}$ distinct molecular structures of interest, we can have up to $n_{\text{max}}=n_{0}(q+1)$ multicules. It is a set of $n\leq n_{\text{max}}$ multicules which we consider here.
Computational duality is a particularly interesting proposition in that is is capable of converting essentially any collection of disparate types of biological processes into a single coherent process type. Abstracting small-molecule structural information and position as part of a collective state can, in principle, be extended to discretizations of other properties such as conformational states of proteins, methylation states of DNA. Versions of this concept appear in systems theory in various contexts, and biophysics allows all properties to be represented as abstract analogs to molecular structures, and all processes as abstract analogs to chemical reactions. The general nature of this view is the reason for the term multicule rather than particule. This combined state space is of course computationally intractably large if one wished to include the entirety of most given biological processes, but reduced state spaces can always be chosen for the variables most relevant to a given scientific question. Perhaps even more interesting is that from a mathematical perspective generalized multicules provide a single, unified framework for all of biology.
RESOURCE AVAILABILITY
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact, Bryce-Allen Bagley (bbagley@stanford.edu).
Materials availability
No experimental materials are associated with this study.
Data and code availability
  • Because this study focused on the mathematical aspects of an applied theory problem in mathematical biology, no code or experimental data were used.
ACKNOWLEDGMENTS
This work was funded by a research grant from the Stanford University Physician-Scientist Training Program. We would like to thank Mykel Kochenderfer, Arec Jamgochian, and Benjamin van Roy for guidance in literature review and recommendations of relevant papers in the fields of POMDPs and RL, and the former two for helpful feedback on conceptualization in the early stages of the project. Additionally, we thank the late Christopher T. Walsh for valuable discussions of underlying concepts in the early stages of the project and for his mentorship. He will be missed.
AUTHOR CONTRIBUTIONS
Conceptualization, B.A.B.; methodology, B.A.B.; investigation, B.A.B., N.K.; writing-–original draft, B.A.B., N.K.; writing-–review & editing, B.A.B. and C.K.P.; funding acquisition, B.A.B. and C.K.P.; resources, B.A.B. and C.K.P.; supervision, C.K.P.
DECLARATION OF INTERESTS
We have no competing interests to declare in relation to this paper.
DECLARATION OF GENERATIVE AI AND AI-ASSISTED TECHNOLOGIES
During the preparation of this work, the authors made no use of Generative AI or AI-assisted technology.
SUPPLEMENTAL INFORMATION INDEX
1: Derivation of the Evolutionary Dynamics Model. This section contains the intermediate steps between the description of the problem statement and that of the time-evolution operator seen in the body of the paper.
2: Model-Predictive Control of the Eco-evolutionary Game. This section describes the Bellman operator and Bellman function, particularly as relates to this paper.
3: Reward Function Estimation. This section explains the method for approximating the integral within the Bellman equation, and details the estimation error and computational complexity properties associated with it.
4: Derivation of Regret Bounds. This section details the calculation of the regret bound for the application of the PSRL reinforcement learning algorithm to the evolutionary game control problem we study in this paper.
5: Statistical Properties of Unlearnable Error $\varepsilon$. This section details the derivation and proof of properties of the component of error which represents a minimum below which the learning algorithm could not improve, including proving that it does not impact the bound on the expectation value of regret.